Session: May 1998 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Cattani Alessandro at Thu May 28 06:37:16 METDST 1998) Success
Esperiment is terminated with good result

Torun (K. Borkowski at Thu May 28 09:30:15 METDST 1998) Success
No known problems.

Westerbork (Tony Foley at Thu May 28 09:59:29 METDST 1998) Minor failures
Problems with decoder, but data looks OK. Readout of weather data stopped at 00:35UT, so all values after that time are the same.

Onsala (J.Conway/A.Polatidis at Thu May 28 15:29:12 METDST 1998) Success
No known problems

Jodrell (Alastair Gunn at Mon Jun 1 15:38:52 METDST 1998) Success
FS time corrected at start of experiment. No known problems.

Cambridge (Alastair Gunn at Mon Jun 1 15:39:28 METDST 1998) Minor failures
Vacuum lost about 01:00 UT, restarted on scan 14R at 02:15 UT. LO contamination in BBCs 2 and 4 USB - all other BBC sidebands look okay. No other known problems.

Effelsberg (dgraham at Tue Jun 2 18:25:41 METDST 1998) Minor failures
No anomalies noted by operator.

Calibration (Jean\-Francois Desmurs at Mon Jun 8 13:15:35 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name gl34.antab. Read the comments in the file to get further information.

Noto (C.Trigilio at Mon Jun 8 17:10:06 METDST 1998) Success